[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Aug 16, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Aug 17 02:50:42 UTC 2016


HSU Shift Log for Tue Aug 16 2016 22:50:35 GMT-0400 (EDT)

Shift: R

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: ASCAT-A data over TD-6/AL982016 did not arrive. The  
swath that included the center started well to the north at 0000 UTC, with  
no data earlier in the orbit. The operational vectors did arrive at FNMOC,  
but ambiguities are not available there. Since the swath starts at 0000  
UTC, could this be related to the change of day? (Note that ASCAT-B files  
do *not* show this problem!)

Other Techincal issues: Call from the SDM to notify us that due to  
unspecificed tech issues the 18Z GFDL and HWRF data would be about 90  
minutes late. However, the I versions seem to be avialble for the 00Z  
guidance.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160817/6b4853ee/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list